# Multimodal Emotion Recognition in Conversations: A Survey of Methods, Trends, Challenges and Prospects

### v1

@@ 6 Challenges and Prospects
With the growing use of interactive machine applications, MERC has become a critical research area. Applications in human-computer interaction Ahmad et al. (2024); Moin et al. (2023), healthcare Ayata et al. (2020); Islam et al. (2024), education Villegas-Ch et al. (2025); Vani and Jayashree (2025), and virtual collaboration demand robust and adaptable emotion recognition technologies that function effectively in naturalistic and dynamic environments. Yang et al. (2022) investigated MER in contexts affected by face occlusions, such as those introduced by surgical and fabric masks. Khan et al. (2024) studied contactless techniques in MER, surveying a range of nonintrusive modalities (e.g., visual cues, physiological signals). Huang (2024) developed a MER system for online learning to enable real-time monitoring and feedback on learners’ emotional states. These studies highlight key directions for advancing MERC systems, particularly to make them more robust and context-aware. The ongoing research should continue to focus on real-world deployment scenarios.
@@ Acknowledgments
The authors acknowledge the use of ChatGPT exclusively to refine the text for grammatical checks in the final manuscript. This work was supported in part by the Guangdong Basic and Applied Basic Research Foundation under Grant 2023A1515011370, the National Natural Science Foundation of China (32371114), the Characteristic Innovation Projects of Guangdong Colleges and Universities (No. 2018KTSCX049), and the Guangdong Provincial Key Laboratory [No.2023B1212060076]. ZG is supported by the National Science Foundation via ARNI (The NSF AI Institute for Artificial and Natural Intelligence), under the Columbia 2025 Research Project ("Towards Safe, Robust, Interpretable Dialogue Agents for Democratized Medical Care").

### v2

@@ 6 Challenges and Prospects
MERC Application. With the growing use of interactive machine applications, MERC has become a critical research area. Applications in human-computer interaction Ahmad et al. (2024); Moin et al. (2023), healthcare Ayata et al. (2020); Islam et al. (2024), education Villegas-Ch et al. (2025); Vani and Jayashree (2025), and virtual collaboration demand robust and adaptable emotion recognition technologies that function effectively in naturalistic and dynamic environments. Yang et al. (2022) investigated MER in contexts affected by face occlusions, such as those introduced by surgical and fabric masks. Khan et al. (2024) studied contactless techniques in MER, surveying a range of nonintrusive modalities (e.g., visual cues, physiological signals). Huang (2024) developed a MER system for online learning to enable real-time monitoring and feedback on learners’ emotional states. These studies highlight key directions for advancing MERC systems, particularly to make them more robust and context-aware. The ongoing research should continue to focus on real-world deployment scenarios. Future MERC systems could benefit from incremental learning techniques and user-in-the-loop feedback mechanisms that allow adaptation in dynamic, real-time environments.
Expanding the Modality Space in Emotion Recognition. Current MER systems mainly rely on vision, audio, and text, given their accessibility and prevalence in datasets. Yet human emotion is expressed through additional channels such as gaze, gestures, posture, and physiological signals (e.g., heart rate, skin conductance, brain activity) Noroozi et al. (2021); Udahemuka et al. (2024); Wang et al. (2023); He et al. (2020); Liu et al. (2024c). These modalities remain underrepresented due to challenges in collecting synchronized, high-quality data and evolving annotation standards Udahemuka et al. (2024); He et al. (2020); Kim and Hong (2024). Expanding beyond the traditional three can reduce reliance on potentially misleading cues and open new application domains. For instance, bio-signals and contextual cues could enhance emotion sensing in health and education, while wearable sensors and eye-trackers may enable emotion-aware experiences in VR, driver monitoring, or social robotics. In summary, gaze, physiological, and other embodied modalities are promising but underexplored in affective computing. Future work should explore how to systematically integrate these diverse modalities into large-scale benchmarks and develop models capable of robustly leveraging them in real-world scenarios.
@@ Acknowledgments
The authors acknowledge the use of ChatGPT exclusively to refine the text for grammatical checks in the final manuscript. This work was supported in part by the Guangdong Basic and Applied Basic Research Foundation under Grant 2023A1515011370, the National Natural Science Foundation of China (32371114), the Characteristic Innovation Projects of Guangdong Colleges and Universities (No. 2018KTSCX049), and the Guangdong Provincial Key Laboratory [No.2023B1212060076]. YL acknowledges support from the Duke Power Endowed Professorship at the School of Business, North Carolina Central University, and the IBM-HBCU Quantum Center Faculty Fellowship. ZG is supported by the National Science Foundation via ARNI (The NSF AI Institute for Artificial and Natural Intelligence), under the Columbia 2025 Research Project ("Towards Safe, Robust, Interpretable Dialogue Agents for Democratized Medical Care").